A rectangular room is 20 ft long, 12 feet wide and 10 ft high. What is the max distance from one corner to another

Answer by checkley79(3341)

YOU HAVE A RIGHT TRIANGLE A^2+B62=C^2
20^2+12^2=C^2
400+144=C^2
544=C^2
C=SQRT544
C=23.3 FEET ANS. FOR THE DIAGONAL.